OpenStudy (anonymous):

Find the mass of a flying discus that has a net force of 1.05 newtons and accelerates at 3.5 m/s2.

OpenStudy (anonymous):

F = m . a m = F / a m = 1.05 / 3.5 m= 0.3

OpenStudy (anonymous):

can i have a medal please:)

OpenStudy (whpalmer4):

if only you'd supplied a unit...